Use synthetic division to find the quotient and the remainder. (x5 + x2 - x?) + (x-2) Q(x) = 0 R(x) =
Answer #1

The solution is as below:

x + x² - x²) + (-2) 2-2 =0 >x=2 2 1 0 1 -1 0 0 24 10 18 36 1 2 5 9 18 (36) Therefore, 8(x) = x + 2x² + 5x² + 9x + 18 R(x) = 3